Jim and his daughter's combined age is 31. Jim is exactly 30 years older than his daughter. How old is his daughter?
Answer by stanbon(75887)

Jim and his daughter's combined age is 31. Jim is exactly 30 years older than his daughter. How old is his daughter?
--------------
Equations:
j + d = 31
j = d + 30
----
Substitute for for "j" and solve for "d":
d+30+d = 31
2d = 1
d = 1/2 (daughter = 6 months)
-----
Solve for "j":
j = d+30
j = (1/2) + 30
j = 30 yrs and 6 months (jim's age now)
===========================================
Cheers,
Stan H.
=========